# That is Unacceptable: the Moral Foundations of Canceling

### v1

@@ Abstract
Soda Marem Lo1,2, Oscar Araque3, Rajesh Sharma4, Marco Antonio Stranisci1,2, 1Università degli Studi di Torino, 2aequa-tech, 3Universitad Complutense de Madrid, 4Tartuk Ülikool Correspondence: marcoantonio.stranisci@unito.it
@@ 1 Introduction
The main objective of our research is to provide the first study of canceling attitudes through the lens of people’s different moral perspectives. To this aim, we developed the Canceling Attitudes DEtection (CADE) dataset: a corpus of canceling incidents gathered from YouTube. The corpus includes six videos regarding controversial events about celebrities and comments that have been annotated for the presence of canceling attitudes against them. Given the social relevance of the task, we involved three types of stakeholders (activists, researchers, and students) in a participatory annotation lab Delgado et al. (2023) during which we constantly received feedback from them about the annotation process and the potential uses of CADE in downstream applications for content moderation. The moral perspectives of annotators are identified according to the Moral Foundations Theory (MFT) Graham et al. (2008), which enables the identification of people’s moral profiles by ranking their attitudes towards five foundations. By developing CADE we investigate two research questions.
RQ1: What is the impact of an individual’s morality in evaluating canceling attitudes? CADE enables a systematic study of human disagreement that is not limited to individuals’ sociodemographics Sap et al. (2019); Frenda et al. (2023) but also considers their moral stance towards canceling attitudes. We clustered annotators on the basis of their moral profiles and observed how they evaluate YouTube comments on controversial events involving celebrities. The analysis shows that individuals’ morality is event-focused: people with different moral profiles tend to evaluate differently the canceling attitudes against specific celebrities, independently of their demographic traits.
@@
This paper is organized as follows. In Section 2 we present relevant related work; in Section 3 the creation of the corpus and in Section 4 its analysis. In Section 5 we present two experiments relative to our two RQs. Section 6 summarizes our findings.
@@ 2 Related Work
The growing interest of the NLP community in moral pluralism Graham et al. (2008); Schwartz (2012) resulted in the creation of resources Araque et al. (2020); Hoover et al. (2020) aimed at analyzing the impact of morality in perceiving harmful contents Stranisci et al. (2021); Davani et al. (2024). Moreover, there is an active line of research that studies the morality of LLMs Abdulhai et al. (2024); Röttger et al. (2024). Our work apply moral pluralism to understand canceling attitudes in social media platforms, where the morality of LLMs can have an impact in content moderation.
@@ 3 The CADE Dataset
In this section we present the process that led to the creation of the Canceling Attitudes DEtection (CADE) dataset: a corpus of YouTube comments annotated for the study of canceling attitudes against controversial events. The section is organized as follows: we first present the data collection (Section 3.1) phase, then we describe the design of the annotation task (Section 3.2), and present the Annotation Lab: the participatory approach that we adopted to recruit annotators (Section 3.3).
@@ 3.1 Data collection
As described in Section 2, cancel culture is characterized by the online public shaming of celebrities for their actions or behaviors. In a NLP perspective, this can be conceived as an event-centric task Chen et al. (2021), where the characteristics of the event orient the interest of the research. Relying on this assumption we chose YouTube as source for our data collection. We leveraged the public APIs to obtain communicative situations where the presence of canceling attitudes can be assessed.
Firstly, we selected controversial events related to six celebrities that refer to various topics: J.K. Rowling (homo-transphobia), Kanye West (antisemitism), Lizzo (harassment), Halle Bailey (anti-woke culture), Ellen DeGeneres (bullying), Andrew Tate (sexual assault). For each of them, we manually selected a video of a news broadcast reporting the event the target celebrity got canceled for. Since the annotators had to watch it before completing the annotation task, we opted for the most viewed video among those that did not exceed 4 minutes. For each video we extracted all the comments and randomly sampled 350 comments for each celebrity, resulting in a total of $2,100$ texts for the annotation.
@@ 3.2 Design of the annotation task
For each of the six target celebrities, we prepared a summary of the controversial event that introduced the topic to the annotators (Appendix B) and then asked them to watch a YouTube video on the controversial event (Appendix C). Following this step, they could annotate the stance and the unacceptability of comments about celebrities. First, annotators must evaluate whether the YouTube user intended to attack, defend or was neutral towards the controversial event involving the target celebrity (stance). Annotators must then evaluate the social unacceptability of the comment, choosing on a scale that ranges from 1 (totally acceptable) to 4 (totally unacceptable).
@@ 3.3 Annotation Lab
The annotation lab is structured in three main steps: i) engagement: we presented a set of slides through a shared video in which we provided a definition of the phenomenon of cancel culture, we shared our research objective and its social impact and explain the whole annotation process; ii) annotation: people were asked to fill out the MFT questionnaire and to annotate YouTube comments according to our annotation scheme; iii) feedback: we asked participants to provide feedback about the annotation process and share their ideas about potential applications of a technology for the identification of canceling attitudes. To this aim, we adopted two methodologies: the administration of a checkout questionnaire, and the organization of two focus groups.
@@ 4 Corpus Analysis
In this section we analyse the CADE corpus. [^2] The repository of the dataset will be made available under a CC BY license after the anonymity period. We first describe the composition of annotators (Section 4.1), then the agreement between them (Section 4.2), finally we describe whether annotators moral profiles correlate with their sociodemographic traits (Section 4.3).
@@ 4.1 Annotation Process
The annotation task involved $57$ annotators belonging to three stakeholders: $30$ students, $17$ activists, and $10$ researchers. In addition, each annotator voluntarily shared their information about gender identity, and age, which was grouped into generations (Boomer, GenX, GenY and GenZ), nationality, ethnicity, education level and employment status. All the sociodemographic data are reported in Table 3 in Appendix A. Having adopted a participatory approach to annotators recruitment (Section 3.3) that focuses on engaging with people who are interested in the phenomenon rather than hiring crowdworkers, the pool of annotators is not balanced along all sociodemographic axes.
Each annotator is asked to annotate a subset of $210$ comments gathered from YouTube (Section 3.1): $35$ for each celebrity and the controversial event related to them. After cleaning unrelated comments, the final corpus includes $2,094$ texts and $11,935$ annotations. Each text has been annotated an average of $5.7$ times (with a median of $6$). Annotators had the option to refrain from assigning a label. We eliminated those who did not finish more than one-third of the task.
@@ 4.2 Inter Annotator Agreement
To assess the Inter Annotator Agreement (IAA) between annotators we employed Krippendorff’s Alpha Krippendorff (2011), which handles both agreement by chance (as the more common Cohen’s Kappa agreement score), and incomplete annotations. We computed the IAA per sample and averaged them, resulting in a moderate agreement on stance ($0.501$), and a fair agreement on acceptability ($0.222$). The strong difference between the two tasks highlights how people tend to agree more when judging YouTube users’ intentions rather than when they are asked to evaluate which messages they consider unacceptable Davani et al. (2024) Labeling the stance expressed by a comment results less subjective than stating whether the comment is to consider unacceptable. We report the scores broken down by sample in Table 5 in Appendix C.
@@ 4.3 Annotators Moral Profiles
As a third part of the analysis of our annotation corpus, we assign annotators’ moral profiles by leveraging the results of the MFT Questionnaire (Section 3.2) and test to which extent they correlate with the following sociodemographic characteristics: gender, age, and stakeholder type. [^3] Since the recruitment of stakeholder has been performed in European countries, we decided not to include nationality and ethnicity as variables of our study, as they result to be unbalanced.We also excluded the education level and employment status because they are strongly dependent on the type of stakeholders involved during the annotation, which is the only social condition monitored by design.
@@
Figure 1: Heatmap visualization for cluster analysis
We performed a qualitative analysis of the obtained moral clusters to understand which patterns emerge from their composition. First, we analyzed whether the clusters were consistent with the scores that individuals obtained on each foundation. The heatmap shown in Figure 1 represents the mean score of the five foundations for each cluster. Looking at the higher divergence, we computed the Pearson coefficient between the clusters and each of the moral foundations, reporting a correlation with loyalty ($r=0.598$ with $p-value=9.191\cdot 10^{-7}$), authority ($r=0.759$ with $p-value=7.562\cdot 10^{-12}$) and purity ($r=0.792$ with $p-value=2.063\cdot 10^{-13}$).
These results are theoretically motivated by the MFT theory, which distinguishes between individual binding foundations (care, harm), rooted on the preservation of individual freedom, and group-binding foundations (loyalty, authority, and purity) which focus on the duties of individuals towards their social groups. In this sense, cluster_1 is characterized by a higher moral attitude towards belonging to a group.
@@
Gender identity Female 23 (67.7%) 11 (32.3%) (
@@
Table 1: Composition of the clusters (CL_0 and CL_1) in respect to annotators’ gender identity, generation and role as stakeholder
Once moral clusters have been validated against the MFT, we investigated again the presence of sociodemographic patterns in clusters to gain more insights from their potential correlation with moralities. Table 1 shows the composition of moral clusters broken down by annotators’ gender, generation, and role as stakeholders. As it can be observed, the distribution of moral clusters along the gender axis is uniform: 32% of women and 20% of men belong to cluster_1, showing a distribution that can be explained by annotators’ gender imbalance. This is even more emphasized if generations are observed: moral clusters are perfectly distributed in Generation Y and Generation Z, despite the age of annotators being highly unbalanced towards the latter. This suggests that, in the context of this research, age is not a factor in determining the morality of annotators. Observing the distribution of moral clusters among stakeholders shows interesting patterns, instead. If on one side 36% of students and 30% of researchers belong to cluster_1, only 11% of activists belong to this cluster.
@@ 5 Experiments
In this section we leverage the CADE dataset to answer the two Research Questions outlined in Section 1. In Section 5.1 we analyze human disagreement in the evaluation of canceling attitudes with a specific focus on the impact of morality in this task. In Section 5.2 we study how LLMs classify canceling attitudes, analyzing whether they follow specific patterns that align them to moral or sociodemographic characteristics of annotators.
@@
5.1 STUDY 1: What is the impact of individual’s morality in evaluating canceling attitudes?
@@
The first experiment aims to analyze the human disagreement of annotators in perceiving canceling attitudes, focusing in particular on the influence of their moral profiles (Section 4.3) in their evaluation. Given the type of dataset, which replicates online situations triggered by controversial events, we analyze to which extent human disagreement is determined by the type of events and celebrities targeted by YouTube comments.
@@
The first step of the study is the analysis of stance and unacceptability emerging from YouTube comments, broken down by celebrity. Table 2 shows the annotations aggregated by celebrities (the columns) and moral and sociodemographic groups (the rows). For each celebrity, the relative distribution of stance labels (a-d-n) and the average unacceptability scores are reported.
@@
The table shows that canceling attitudes vary significantly among celebrities. Looking at the last row, only a few comments attack J.K. Rowling ($15\%$) for her transphobic declarations and Andrew Tate ($14\%$) for the accusation against him of sexual assault. Conversely, Lizzo suffers a high number of attacks ($64\%$) for her misconduct against a member of her staff, as well as Ellen DeGeneres ($67\%$) for her bullying attitudes in her working environment. The perceived unacceptability of YouTube comments appears to be orthogonal to stance, since comments can be perceived as highly unacceptable regardless they attack celebrities or not. For instance, J.K. Rowling is mostly defended but the unacceptability score is high. This means that there are many comments that defend her by adopting canceling attitudes against other targets: in these cases Daniel Radcliffe and Emma Watson, who called out against her transphobic claims. Lizzo is mostly attacked and comments about her scored the highest unacceptability ($2.6$): this means that she is the actual target of canceling attitudes. Comments on Andrew Tate and Kanye West, who mainly defend them, also score the lowest unacceptability scores ($1.7$), showing that their focus is more on defending them rather than attacking other targets.
@@
The analysis of annotations broken down by moral cluster, gender, and stakeholder shows lower variation among moral and sociodemographic axes than one emerging from the analysis of celebrities (Table 2, last column). The percentage of comments annotated as attacking celebrities is uniform along morality, gender, and type of stakeholder, ranging from $0.36$ to $0.38$. The percentage of comments evaluated as defending celebrities shows more variation along moral clusters (cluster_0: $42\%$ versus cluster_1: $37\%$) and stakeholders (activists: $44\%$ versus students: $38\%$). Variation in the evaluation of unacceptability is even less significant. Annotators’ moral profiles appear to have no impact on the perception of this phenomenon: the average unacceptability is $2.03$ for both groups. Stakeholders show some variation, with the average unacceptability provided by activists ($2.09$) that diverges by $0.07$ points from the average unacceptability provided by students ($2.02$). Differences along the gender axis are more significant, as it is possible to observe a high variation between the unacceptability score assigned by women ($2.07$) with one assigned by men ($1.96$).
@@
For each axis of interest (e.g., morality) we computed the relative distribution of stance and the average unacceptability scores and computed the Chi-squared test between the distribution of labels assigned by annotators belonging to different groups (e.g., cluster_0 vs cluster_1) in order to assess if there is a significant divergence between their evaluations. All the cases where the Chi-squared test shows a statistically significant divergence ($p<0.05$) are highlighted in bold in Table 2. As it can be observed in the table, the different perception of stance among annotators with different moral profiles clearly emerges if specific celebrities are examined. The two moral clusters diverge in the evaluation of attacking comments against J.K. Rowling (cluster_0: $13\%$ versus cluster_1: $21\%$) and Kanye West (cluster_0: $20\%$ versus cluster_1: $27\%$) and of defending comments against Andrew Tate (cluster_0: $71\%$ versus cluster_1: $61\%$). Similar variations can be observed at the level of the gender axis and stakeholder axis. Women and men diverge the most in evaluating the defending stance on comments on Andrew Tate (men: $74\%$ versus women $66\%$); students ($18\%$) perceive more attacking comments against J.K. Rowling than activists ($12\%$) and researchers ($12\%$).

### v3

@@ Abstract
Soda Marem Lo1,2, Oscar Araque3, Rajesh Sharma4,5, Marco Antonio Stranisci1,2, 1Università degli Studi di Torino, 2aequa-tech, 3Universidad Politécnica de Madrid, 4University of Tartu, Estonia, 5Plaksha University, India Correspondence: marcoantonio.stranisci@unito.it sodamarem.lo@unito.it
@@ 1 Introduction
The main objective of our research is to provide the first study of canceling attitudes through the lens of people’s different moral perspectives. To this aim, we developed the Canceling Attitudes DEtection (CADE) dataset: a corpus of canceling incidents gathered from YouTube. The corpus includes six videos regarding controversial events about celebrities and comments that have been annotated for the presence of canceling attitudes against them. Given the social relevance of the task, we involved three types of stakeholders (activists, researchers, and students) in a participatory annotation lab Delgado et al. (2023). Throughout the lab, we constantly received feedback from them about the annotation process and the potential uses of CADE in downstream applications for content moderation. The moral perspectives of annotators are identified according to the Moral Foundations Theory (MFT) Graham et al. (2008), which enables the identification of people’s moral profiles by ranking their attitudes towards five foundations: care of the most vulnerable members of the community, fairness in the cooperation with others, loyalty to the group, respect of the authority, and sense of purity. By developing CADE we investigate two research questions.
RQ1: What is the impact of an individuals’ morality in evaluating canceling attitudes? CADE enables a systematic study of human disagreement that is not limited to individuals’ sociodemographics Sap et al. (2019); Frenda et al. (2023) but also considers their moral stance towards canceling attitudes. We clustered annotators on the basis of their moral profiles and observed how they evaluate YouTube comments on controversial events involving celebrities. The analysis shows that individuals’ morality is event-focused: people with different moral profiles tend to evaluate differently the canceling attitudes against specific celebrities, independently of their demographic traits.
@@
This paper is organized as follows. In Section˜2 we present relevant related work; in Section˜3 the creation of the corpus and in Section˜4 its analysis. In Section˜5 we present two experiments relative to our two RQs. Section˜6 summarizes our findings.
@@ 2 Related Work
The growing interest of the NLP community in moral pluralism Graham et al. (2008); Schwartz (2012) resulted in the creation of resources Araque et al. (2020); Hoover et al. (2020) aimed at analyzing the impact of morality in perceiving harmful contents Stranisci et al. (2021); Davani et al. (2024). Moreover, there is an active line of research that studies the morality of LLMs Abdulhai et al. (2024); Röttger et al. (2024). Our work applies moral pluralism to understand canceling attitudes in social media platforms, where the morality of LLMs can have an impact in content moderation.
@@ 3 The CADE Dataset
In this section we present the process that led to the creation of the Canceling Attitudes DEtection (CADE) dataset: a corpus of YouTube comments annotated for the study of canceling attitudes against controversial events. The section is organized as follows: we first present the data collection phase (Section˜3.1), then we describe the design of the annotation task (Section˜3.2), and present the Annotation Lab: the participatory approach that we adopted to recruit annotators (Section˜3.3).
@@ 3.1 Data collection
As described in Section˜2, cancel culture is characterized by the online public shaming of celebrities for their actions or behaviors. In a NLP perspective, this can be conceived as an event-centric task Chen et al. (2021), where the characteristics of the event orient the interest of the research. Relying on this assumption we chose YouTube as source for our data collection. We leveraged the public APIs [^1] Besides the comment, we requested the username, immediately anonymised, the date and time when the comment was posted, and the number of likes. However, the final corpus does not contain any of the collected metadata, which will not be released. to obtain communicative situations where the presence of canceling attitudes can be assessed.
Firstly, we selected controversial events related to six celebrities that refer to various topics: J.K. Rowling (homo-transphobia), Kanye West (antisemitism), Lizzo (harassment), Halle Bailey (anti-woke culture), Ellen DeGeneres (bullying), Andrew Tate (sexual assault). For each of them, we manually selected a video of a news broadcast reporting the event the target celebrity got canceled for. We limited potential biases linked to the political orientation of the source by balancing between right- and left-leaning channels, as reported in Table˜5 (Appendix C). Since the annotators had to watch the video before completing the annotation task, we opted for the most viewed clip among those that did not exceed 4 minutes. For each video we extracted all the comments and randomly sampled 350 comments for each celebrity, resulting in a total of $2,100$ texts for the annotation.
@@ 3.2 Design of the annotation task
For each of the six target celebrities, we prepared a summary of the controversial event that introduced the topic to the annotators (Appendix B) and then asked them to watch a YouTube video on the controversial event (Appendix C). Following this step, they could annotate the stance and the unacceptability of comments about celebrities. First, annotators had to evaluate whether the YouTube user intended to attack, defend or was neutral towards the controversial event involving the target celebrity (stance). Then annotators had to evaluate the social unacceptability of the comment, choosing on a scale that ranges from 1 (totally acceptable) to 4 (totally unacceptable).
@@ 3.3 Annotation Lab
The annotation lab is structured in three main steps: i) engagement: we presented a set of slides through a shared video in which we provided a definition of the phenomenon of cancel culture, we shared our research objective and its social impact and explain the whole annotation process; ii) annotation: people were asked to fill out the MFT questionnaire and to annotate YouTube comments according to our annotation scheme; iii) feedback: we asked participants to provide feedback about the annotation process and share their ideas about potential applications of a technology for the identification of canceling attitudes. Differently from traditional data annotation methodology in the NLP field, we preferred giving priority to the direct participation of the annotators by stating the purpose of the study. We wanted to make them aware of the goals of the research they contributed to, giving space to the possibility of discussing and explicitly sharing these objectives. To this aim, we adopted two methodologies: the administration of a checkout questionnaire, and the organization of two focus groups.
@@ 4 Corpus Analysis
In this section we analyse the CADE corpus. [^3] https://github.com/aequa-tech/canceling-attitudes We first describe the composition of annotators (Section˜4.1), then the agreement between them (Section˜4.2), finally we describe whether annotators moral profiles correlate with their sociodemographic traits (Section˜4.3).
@@ 4.1 Annotation Process
The annotation task involved $57$ annotators belonging to three stakeholders: $30$ students, $17$ activists, and $10$ researchers. In addition, each annotator voluntarily shared their information about gender identity, and age, which was grouped into generations (Boomer, GenX, GenY and GenZ), nationality, ethnicity, education level and employment status. All the sociodemographic data are reported in Table˜4 in Appendix A. Having adopted a participatory approach to annotators recruitment (Section˜3.3) that focuses on engaging with people who are interested in the phenomenon rather than hiring crowdworkers, the pool of annotators is not balanced along all sociodemographic axes.
Each annotator is asked to annotate a subset of $210$ comments gathered from YouTube (Section˜3.1): $35$ for each celebrity and the controversial event related to them. After cleaning unrelated comments, the final corpus includes $2,094$ texts and $11,935$ annotations. Each text has been annotated an average of $5.7$ times (with a median of $6$). Annotators had the option to refrain from assigning a label. We eliminated those who did not finish more than one-third of the task. Appendix E reports more detailed statistics on the annotations, and Appendix F textual examples.
@@ 4.2 Inter Annotator Agreement
To assess the Inter Annotator Agreement (IAA) between annotators we employed Krippendorff’s Alpha Krippendorff (2011), which handles both agreement by chance (as the more common Cohen’s Kappa agreement score), and incomplete annotations. We computed the IAA per sample and averaged them. Adopting Landis and Koch’s terminology (1977), it resulted in a moderate agreement on stance ($0.501$), and a fair agreement on acceptability ($0.222$). These results are comparable with other datasets released in a disaggregated fashion on analogous tasks, such as offensive speech detection Sachdeva et al. (2022). The strong difference between stance and acceptability highlights how people tend to agree more when judging YouTube users’ intentions rather than when they are asked to evaluate which messages they consider unacceptable Davani et al. (2024). Labeling the stance expressed by a comment results less subjective than stating whether the comment is to consider unacceptable. We report the scores broken down by sample in Table˜6 in Appendix C.
@@ 4.3 Annotators Moral Profiles
As a third part of the analysis of our annotation corpus, we assign annotators’ moral profiles by leveraging the results of the MFT Questionnaire (Section˜3.2) and testing to which extent they correlate with the following sociodemographic characteristics: gender, age, and stakeholder type. [^4] Since the recruitment of stakeholder has been performed in European countries, we decided not to include nationality and ethnicity as variables of our study, as they result to be unbalanced.We also excluded the education level and employment status because they are strongly dependent on the type of stakeholders involved during the annotation, which is the only social condition monitored by design.
@@
Figure 1: Heatmap of the mean scores for the five foundations across clusters.
We performed a qualitative analysis of the obtained moral clusters to understand which patterns emerge from their composition. First, we analyzed whether the clusters were consistent with the scores that individuals obtained on each foundation. The heatmap shown in Figure˜1 represents the mean score of the five foundations for each cluster. Looking at the higher divergence, we computed the Pearson coefficient between the clusters and each of the moral foundations, reporting a correlation with loyalty ($r=0.598$ with $p-value=9.191\cdot 10^{-7}$), authority ($r=0.759$ with $p-value=7.562\cdot 10^{-12}$) and purity ($r=0.792$ with $p-value=2.063\cdot 10^{-13}$).
These results are theoretically motivated by the MFT theory, which distinguishes between individual binding foundations (care, harm), rooted on the preservation of individual freedom, and group-binding foundations (loyalty, authority, and purity) which focus on the duties of individuals towards their social groups Haidt and Graham (2007). In this sense, cluster_1 is characterized by a higher moral attitude towards belonging to a group.
@@
Gender identity Female 23 (67.7%) 11 (32.3%)
@@
Table 1: Composition of the clusters (CL_0 and CL_1) in respect to annotators’ gender identity, generation and role as stakeholder.
Once moral clusters have been validated against the MFT, we investigated again the presence of sociodemographic patterns in clusters to gain more insights from their potential correlation with moralities. Table˜1 shows the composition of moral clusters broken down by annotators’ gender, generation, and role as stakeholders. As it can be observed, the distribution of moral clusters along the gender axis is uniform: 32% of women and 20% of men belong to cluster_1, showing a distribution that can be explained by annotators’ gender imbalance. This is even more emphasized if generations are observed: moral clusters have the same percentage of Generation Y and Generation Z annotators, despite their age being highly unbalanced towards the latter. This suggests that, in the context of this research, age is not a factor in determining the morality of annotators. Observing the distribution of moral clusters among stakeholders shows interesting patterns, instead. If on one side 36% of students and 30% of researchers belong to cluster_1, only 11% of activists belong to this cluster.
@@ 5 Experiments
In this section we leverage the CADE dataset to answer the two Research Questions outlined in Section˜1. In Section˜5.1 we analyze human disagreement in the evaluation of canceling attitudes with a specific focus on the impact of morality in this task. In Section˜5.2 we study how LLMs classify canceling attitudes, analyzing whether they follow specific patterns that align them to moral or sociodemographic characteristics of annotators.
@@
5.1 STUDY 1: What is the impact of individuals’ morality in evaluating canceling attitudes?
@@
The first experiment aims to analyze the human disagreement of annotators in perceiving canceling attitudes, focusing in particular on the influence of their moral profiles (Section˜4.3) in their evaluation. Given the type of dataset, which replicates online situations triggered by controversial events, we analyze to which extent human disagreement is determined by the type of events and celebrities targeted by YouTube comments.
@@
The first step of the study is the analysis of stance and unacceptability emerging from YouTube comments, broken down by celebrity. Table˜2 shows the annotations aggregated by celebrities (the columns) and moral and sociodemographic groups (the rows). For each celebrity, the relative distribution of stance labels (a-d-n) and the average unacceptability scores are reported.
@@
The table shows that canceling attitudes vary significantly among celebrities. Looking at the last row, only a few comments attack J.K. Rowling ($15\%$) for her transphobic declarations and Andrew Tate ($14\%$) for the accusation against him of sexual assault. Conversely, Lizzo suffers a high number of attacks ($64\%$) for her misconduct against a member of her staff, as well as Ellen DeGeneres ($67\%$) for her bullying attitudes in her working environment. Observing the correlation between the level of unacceptability and stance (Table˜3), it emerges the tendency for attacking to correspond with increased unacceptability, especially compared to a neutral stance. Defend presents more varied results, with a notable percentage of cases showing moderate levels of unacceptability (2 and 3). However, when looking at the celebrity level (Table˜2), the perceived unacceptability of YouTube comments appears to be orthogonal to stance. Comments can be perceived as highly unacceptable regardless they attack celebrities or not. For instance, J.K. Rowling is mostly defended but the unacceptability score is high. Since annotators were asked how much a comment contributed to shaming a target, we infer that such a result suggests many comments defend her by adopting canceling attitudes against others: in these cases Daniel Radcliffe and Emma Watson, who called out against her transphobic claims. Lizzo is mostly attacked and comments about her scored the highest unacceptability ($2.6$): this means that she is the actual target of canceling attitudes. Comments on Andrew Tate and Kanye West, who mainly defend them, also score the lowest unacceptability scores ($1.7$), showing that their focus is more on defending them rather than attacking other targets.
Unacc. Stance % Annotations
1 Attack 15.97
Defend 52.80
Neutral 31.23
2 Attack 41.46
Defend 39.36
Neutral 19,18
3 Attack 52.36
Defend 35.43
Neutral 12.21
4 Attack 74.61
Defend 18.35
Neutral 7.04
Table 3: Distribution of texts labeled as Attack, Defend, and Neutral, broken down by the range of unacceptability. The final column reports the percentage of texts for each stance within each acceptability level, relative to the total number of texts at that acceptability level. As annotators could abstain from labeling, results are computed based on the instances with completed annotations for both tasks, thus excluding $350$ annotations.
@@
The analysis of annotations broken down by moral cluster, gender, and stakeholder shows lower variation among moral and sociodemographic axes than one emerging from the analysis of celebrities (Table˜2, last column). The percentage of comments annotated as attacking celebrities is uniform along morality, gender, and type of stakeholder, ranging from $0.36$ to $0.38$. The percentage of comments evaluated as defending celebrities shows more variation along moral clusters (cluster_0: $42\%$ versus cluster_1: $37\%$) and stakeholders (activists: $44\%$ versus students: $38\%$). Variation in the evaluation of unacceptability is even less significant. Annotators’ moral profiles appear to have no impact on the perception of this phenomenon: the average unacceptability is $2.03$ for both groups. Stakeholders show some variation, with the average unacceptability provided by activists ($2.09$) that diverges by $0.07$ points from the average unacceptability provided by students ($2.02$). Differences along the gender axis are more significant, as it is possible to observe through the variation between the unacceptability score assigned by women ($2.07$) and the one assigned by men ($1.96$).
@@
For each axis of interest (e.g., morality) we computed the relative distribution of stance and the average unacceptability scores and computed the Chi-squared test between the distribution of labels assigned by annotators belonging to different groups (e.g., cluster_0 vs cluster_1) in order to assess if there is a significant divergence between their evaluations. All the cases where the Chi-squared test shows a statistically significant divergence ($p<0.05$) are highlighted in bold in Table˜2. As it can be observed in the table, the different perception of stance among annotators with different moral profiles clearly emerges if specific celebrities are examined. The two moral clusters diverge in the evaluation of attacking comments against J.K. Rowling (cluster_0: $13\%$ versus cluster_1: $21\%$) and Kanye West (cluster_0: $20\%$ versus cluster_1: $27\%$) and of defending comments against Andrew Tate (cluster_0: $71\%$ versus cluster_1: $61\%$). Similar variations can be observed at the level of the gender axis and stakeholder axis. Women and men diverge the most in evaluating the defending stance on comments on Andrew Tate (men: $74\%$ versus women $66\%$); students ($18\%$) perceive more attacking comments against J.K. Rowling than activists ($12\%$) and researchers ($12\%$).